A 12 ohm resistor is in series with an 8 ohm load. What is the total source voltage if the voltage across the 12 ohm resistor is 36 Volts?

1
Expert's answer
2022-05-10T13:21:25-0400

The current through the resistors

"I=\frac{V_1}{R_1}=\frac{36}{12}=3\:\rm A"

The voltage drop across second resistor

"V_2=IR_2=3*8=24\:\rm V"

The total source voltage

"V=V_1+V_2=36+24=60\:\rm V"